q-ROOK POLYNOMIALS AND MATRICES OVER FINITE FIELDS

JAMES HAGLUND

Abstract. Connections between q-rook polynomials and matrices over finite fields are exploited to derive a new statistic for Garsia and Remmel’s q-hit polynomial. Both this new statistic mat and another statistic for the q-hit polynomial ξ recently introduced by Dworkin are shown to induce different multiset Mahonian permutation statistics for any Ferrers board. In addition, for the triangular boards they are shown to generate different families of Euler-Mahonian statistics. For these boards the ξ family includes Denert’s statistic den, and gives a new proof of Foata and Zeilberger’s Theorem that (exc, den) is jointly distributed with (des, maj). The mat family appears to be new. A proof is also given that the q-hit polynomials are symmetric and unimodal.